The average cholesterol content of a certain can goods is 215 milligrams and the cholesterol deviation is 15 milligrams.assume the variable is normally distributed


1
Expert's answer
2021-02-24T12:39:17-0500

Let "X=" the average cholesterol content of a certain can goods in milligrams

"X\sim N(\mu, \sigma^2/n)"

Then "Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0. 1)"

Given "\mu=215mg, \sigma=15mg"


a.  If a single egg is selected, find the probability that the cholesterol content will be more than 220 milligrams.

"n=1"


"P(X>220)=1-P(X\leq220)"

"=1-P(Z\leq\dfrac{220-215}{15/\sqrt{1}})\approx1-P(Z\leq0.33333)"

"\approx0.3694"

b. If a sample of 25 eggs is selected, find the probability that the mean of the sample will be larger than 220 milligrams.

"n=25"


"P(X>220)=1-P(X\leq220)"

"=1-P(Z\leq\dfrac{220-215}{15/\sqrt{25}})\approx1-P(Z\leq1.66667)"

"\approx0.0478"